OpenStudy (anonymous):

For the function y = f(x), what is the ordered pair for the point on the graph when x = b - 2? A. (b - 2, f(b -2)) B. (x, f(b)) C. (x, b - 2) D. (b - 2, f(b)) @HectorH

OpenStudy (biohazard9064):

A.(b-2,f(b-2)

OpenStudy (anonymous):

you sure?

OpenStudy (biohazard9064):

because y=f(x) x=b-2=y= f(b-2)

OpenStudy (biohazard9064):

So it would be A

OpenStudy (anonymous):

ok thnx :)

OpenStudy (biohazard9064):

Np